1/4

Christian Morgenstern, Christian Morgenstern, Christian Morgenstern, Christian Morgenstern,

Christian Morgenstern,

Le loup-garou La lupfantomo The Banshee (An Approach)

The Werewolf

Der Werwolf

tradukita de R. Platteau tradukita de Bertilo Wennergren tradukita de Max Knight

tradukita de Alexander Gross

Un loup-garou, une certaine nuit, Laissa sa moitié et puis se rendit Auprès de la tombe d’un maître

d’école Et lui dit, “ sur moi dis tout, c’est

ton rôle ! ”

Jen lupfantom’ de fa-

mili’ meznokte iam kuris

for al tombo de la profe-

sor’. kaj petis pri konjuga-

ci’.

One night, a banshee slunk

away from mate and child, and in

the gloom went to a village teacher’s

tomb, requesting him: “Inflect me,

pray.”

A Werewolf, troubled by his na-

me, Left wife and brood one night

and came To a hidden graveyard to enlist The aid of a long-dead philolo-

gist.

Ein Werwolf eines Nachts

entwich von Weib und Kind und

sich begab an eines Dorfschullehrers

Grab und bat ihn: �Bitte, beuge

mich!“

Un loup-garou, une certaine nuit, Laissa sa moitié et puis se rendit Auprès de la tombe d’un maître

d’école Et lui dit, “ sur moi dis tout, c’est

ton rôle ! ”

Do diktis tiu dokta ul’ starante sur la kructa-

bul’ dum lup’ senmova kiel

splinto atentis vortojn de l’

mortinto.

The village teacher climbed

up straight upon his grave stone with

its plate and to the apparition said who meekly knelt before the

dead:

“Oh sage, wake up, please don’t

berate me,” He howled sadly, “Just conjuga-

te me.” The seer arose a bit unsteady Yawned twice, wheezed once, and then was ready.

Der Dorfschulmeister

stieg hinauf auf seines Blechschilds Messingknauf und sprach zum Wolf, der

seine Pfoten geduldig kreuzte vor dem

Toten:

…  …  …  …  …

2/4

“ - On peut à ton sujet puisque nous

y sommes Se demander encor bien des choses

en somme. Loup-garoù ? mais avant Il faut se dire : loup-garquand ?

De “lupfantomo”, sim-

pla rekta, tra “lupfintomo”, plej

perfekta, plu “lupfontomo”, iom

rara, ĝis “lupfuntomo”, ho,

koŝmara.

“The banSHEE, in the sub-

ject’s place; the banHERS, the possessi-

ve case. The banHER, next, is what

they call objective case–and that is

all.”

“Well, ‘Werewolf’ is your plural

past, While ‘Waswolf’ is singularly

cast: There’s ‘Amwolf’ too, the pre-

sent tense, And ‘Iswolf,’ ‘Arewolf’ in this

same sense.”

�Der Werwolf“ – sprach der gute Mann, �des Werwolfs, Genitiv so-

dann, dem Werwolf, Dativ, wie

man’s nennt, den Werwolf, - damit hat’s ein End!“

Loup-garcomment ? et puis Bien évidemment loup-garqui ? ” Toutes ces questions semblèrent au

loup bien absconses, Car en fait ce qu’il voulait c’était les

réponses.

Ĉi participoj plaĉis

multe al la fantomo, kaj re-

zulte ĝi petis: “Legu plu el

PIV-o pri participoj en pasi-

vo.”

The banshee marveled at

the cases and writhed with pleasure,

making faces, but said: “You did not add,

so far, the plural to the singular!”

“I know that–I’m no mental

cripple– The future form and participle Are what I crave,” the beast re-

plied. The scholar paused–again he

tried:

Dem Werwolf schmeichel-

ten die Fälle er rollte seine Augenbälle. �Indessen“, bat er, �füge

doch zur Einzahl auch die Mehrzahl noch!“

“ - Point de réponses, dit le puits de

savoir, Etre imaginaire, nul ne peut te voir

! ” A ces mots l’animal éclata tout en

larmes. “ - Oh!, même les baisers de ma dou-

ce moitié La viro ĉagreniĝis ta-

men: “Jam fermis mi per fin’ kaj amen, ĉar netransiraj bestoj

feblas kaj la pasiv’ neniam

eblas.”

The teacher, though, admit-

ted then that this was not within his

ken. “While bans are frequent”,

he advised, “A she cannot be plurized.”

“A ‘Will-be-wolf?’ It’s just too

long: ‘Shall-be-wolf?’ ‘Has-been-wolf?’ Utterly wrong! Such words are wounds beyond

all suture– I’m sorry, but you have no fu-

ture.”

Der Dorfschullehrer aber

musste gestehn, dass er von ihr nichts wusste. Zwar Wölfe gab’s in großer Schar,

doch �Wer“ gab’s nur im Sinular.

…  …  …  …  …

3/4

Ne pourrons plus me faire me sentir

tout entier. Rester un loup-garou est un sort qui

m’alarme. ” Mais comme le loup était bien poli Il remercia et puis repartit.

Sed lupfantomo pri

transiro ja spertas pli ol eĉ

vampiro. La lup’ revenis larmo-

kula al sia famili’ ulula.

The banshee, rising clammi-

ly, wailed: “What about my fa-

mily?” Then, being not a learned

creature, said humbly “Thanks” and

left the teacher.

The Werewolf knew better–his

sons still slept At home, and homewards now

he crept, Happy, humble, without apolo-

gy For such folly of philology.

Der Wolf erhob sich tränenblind – er hatte ja doch Weib und

Kind! Doch da er keine Gelehr-

ter eben, so schied er dankend und

ergeben.

…  …  …  …  …

4/4

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Francan de R. Platteau.

Arg-825-1655 (2013-02-13 21:56:38)

Tiu ĉi francigo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en Esperanton de Bertilo Wennergren.

Arg-825-1653 (2013-02-14 23:29:47)

Vidu  la  retejon

http: // bertilow. com/ literaturo/ lupfantomo. html .

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Anglan de Max Knight (civila nomo: Max Kühnel, ∗1909-06-08 – †1993-08-31).

Arg-825-1660 (2013-02-13 22:36:06)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html . Pri la tradukinto Max Knight vidu: http: // www. jbeilharz. de/ morgenstern/ morgenstern_ poems. html kaj http: // www. 123people. de/ ext/ frm? ti= ersonensuche% 20telefonbuch&search_ term= m x% 20k% C3% BChnel&search_country= DE&st= suc e% 20nach% 20personen&target_url= aHR0cDovL d3dy5saXRlcmF0dXJoYXVzLmF0L2luZGV4LnBocD9pZD02ODc4 .

Traduko de la Germana poemo “Der Werwolf” de Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) en la Anglan de Alexander Gross.

Arg-825-1657 (2013-02-13 18:39:45)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .

Verkinto de tiu ĉi Germana poemo estas Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31).

Arg-825-1652 (2013-02-11 18:44:21)